whats the slope of -5,10 and 0,13? sorry i can't put the parenthesis on this
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

whats the slope of (-5,10) and (0,13)?
-----------
slope = (13-10)/(0--5) = 3/5
========
Cheers,
Stan H.
===========

Answer by solver91311(24713)



Use the slope formula:



where and are the coordinates of the given points.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism